OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible. 2the square root of 2. 12the square root of 2. 2the square root of 5. 12the square root of 5.

OpenStudy (anonymous):

\[\sqrt{180} - \sqrt{125}+ \sqrt{5}\] \[2\sqrt{2}\] \[12\sqrt{2}\] \[2\sqrt{5}\] \[12\sqrt{5}\]

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

@mathstudent55

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HI!!

OpenStudy (misty1212):

\[\sqrt{180} - \sqrt{125}+ \sqrt{5}\] right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

if there is anything we can do with this, the first two terms have to be written with some number times \(\sqrt5\) and then we can combine like terms

OpenStudy (misty1212):

for example \[180=36\times 5\] so \[\sqrt{180}=\sqrt{36}\sqrt5=6\sqrt5\]

OpenStudy (misty1212):

try it with \(\sqrt{125}\) and see what you get

OpenStudy (anonymous):

okay so I would do what?

OpenStudy (misty1212):

divide \(125\) by \(5\) and see what you get bet it is a "perfect square"

OpenStudy (anonymous):

25

OpenStudy (misty1212):

yeah and what is \(\sqrt{25}\)?

OpenStudy (anonymous):

5 ?

OpenStudy (misty1212):

lol no not 5? but rather 5!

OpenStudy (misty1212):

so you have \[6\sqrt5+5\sqrt5+\sqrt5\] add like comining like terms

OpenStudy (anonymous):

okay so that would leave us with 6 though

OpenStudy (misty1212):

oh no what is \[6+5+1\]?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

so it would be \[12\sqrt{5}\] right?

OpenStudy (misty1212):

so \[6x+5x+x=12x\] and \[6\sqrt5+5\sqrt5+\sqrt5=12\sqrt 5\]

OpenStudy (misty1212):

yay \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

thanks!